OpenStudy (anonymous):

Does the equation 2x + 8= y represent a direct variation? If so, what is the constant of variation?

OpenStudy (ryuga):

If it is direct variation, you should be able to multiply x by something to get y---no addition or subtraction. y = 3x y = .25x y = (1/3)x these are all direct variation y = 3x + 2 y = .25x - 1 y = (1/3)x + 7 these are not direct variation problems 1-3: direct variation or not? solve each for y and see if there is any addition or subtraction over with the x. If there is, it is NOT direct variation. If you only have something multiplying the x (even an implied 1), the constant will be the coefficient of x. problems 4-12: figure out what you have to multiply x by to get y. In the end you need to get to: y = ?x --where ? is whatever you'll multiply with x to get y. or you could solve for ?--- y = ?x y/x = ?x/x y/x = ? so to find out what ? is, simply divide y by x. Then write your equation in y = ?x format. Hope that gives you enough help to do the work. It shouldn't be too bad.